OpenStudy (anonymous):

Divide. Assume the denominator does not equal zero. 12x^2y-6xy+4x^2/2xy

OpenStudy (luigi0210):

You separate it into two: \(\Large \frac{12x^2y}{2xy}-\frac{6xy}{2xy}+\frac{4x^2}{2xy}\) Now simplify individually.

OpenStudy (anonymous):

how do you simplify it?

OpenStudy (luigi0210):

It can be separated farther if you need it: \[\LARGE \frac{12x^2y}{2xy} =\frac{12}{2}*\frac{x^2}{x}*\frac{y}{y}\] Do you know how to divide variables and powers?

OpenStudy (anonymous):

no :/ im new at this stuff..

OpenStudy (luigi0210):

Rule being, say we have something like \(\Large \frac{x^5}{x^2}\) we simplify by subtratcing exponents like such \(\Large x^{5-2}\), leaving us with \(\Large \frac{x^5}{x^2}=x^3\). Make sense?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so 12x^2y/2xy would be 12xy/y? or just 12x?

OpenStudy (luigi0210):

Close, you forgot that 2 under 12, just do regular division with that. And the y/y cancel leaving a 1. So it'll just be 6x.